OpenStudy (anonymous):

This confuses me. Can you help for a medal? What is the simplified expression for 5^-3 • 2^2 • 5^5 over 2^3 • 5^2 2^2 over 5^2 5 over 2^2 5 over 2 5^2 over 2

OpenStudy (anonymous):

Wouldn't it be C?

OpenStudy (anonymous):

I'm thinking it would be C or D? Can someone please confirm

jimthompson5910 (jim_thompson5910):

There seems to be a typo. I'm getting a different answer which isn't in the list of choices.

jimthompson5910 (jim_thompson5910):

5^(-3)*2^2*5^5 = 100 2^3*5^2 = 200 100/200 = 1/2

OpenStudy (anonymous):

But I don't have that as any choices.:/

jimthompson5910 (jim_thompson5910):

you're absolutely sure the problem is 5^-3 • 2^2 • 5^5 over 2^3 • 5^2 ? can you post a screenshot?

OpenStudy (anonymous):

how do you do that

OpenStudy (anonymous):

I have no way to

jimthompson5910 (jim_thompson5910):

go here http://www.take-a-screenshot.org/ it's a website that has instructions on how to take a screenshot (it varies for different computers)

OpenStudy (anonymous):

Otai one sec ty

OpenStudy (anonymous):

I can't upload anything I only have the option to post a comment

OpenStudy (anonymous):

But they want the simplified fraction, not fully solved

jimthompson5910 (jim_thompson5910):

there's no attach file button below the text box?

OpenStudy (anonymous):

That's why I think it would be c

OpenStudy (anonymous):

No

jimthompson5910 (jim_thompson5910):

can you upload the pic to something like imgur? then host the link

OpenStudy (anonymous):

Hangon

OpenStudy (anonymous):

http://i.imgur.com/IYiK3kh.jpg

OpenStudy (anonymous):

There's the direct link

jimthompson5910 (jim_thompson5910):

thanks

OpenStudy (anonymous):

No problem

jimthompson5910 (jim_thompson5910):

ok so yeah, it looks like \[\Large \frac{5^{-3}\cdot 2^2 \cdot 5^5}{2^3 \cdot 5^2}\] but the problem is that none of the answer choices come up This is what I get \[\Large \frac{5^{-3}\cdot 2^2 \cdot 5^5}{2^3 \cdot 5^2}\] \[\Large \frac{5^{-3+5}\cdot 2^2}{2^3 \cdot 5^2}\] \[\Large \frac{5^{2}\cdot 2^2}{2^3 \cdot 5^2}\] \[\Large \frac{2^2}{2^3}\] \[\Large \frac{4}{8}\] \[\Large \frac{1}{2}\] So strange how 1/2 isn't in the list

OpenStudy (anonymous):

That's what confuses me on why I'm not sure /:

jimthompson5910 (jim_thompson5910):

this is just one of those problems where they messed up big time

OpenStudy (anonymous):

So what one should I put down

jimthompson5910 (jim_thompson5910):

I'm not sure. Any seem likely as to what they meant to say. You'll have to ask your teacher about it since s/he will know what to do. Most likely, the teacher will give you free points back for this one since it's unfair

OpenStudy (anonymous):

Okay thank you

jimthompson5910 (jim_thompson5910):

no problem